OpenStudy (dumbsearch2):

A company that sells computer keyboards has a revenue of 200x - 2x2 dollars and a profit of 3x2 + 245x + 210 for every keyboard x. If Profit = Revenue Cost, find the companys cost equation.

OpenStudy (dumbsearch2):

Thank you SOOO much for your help... :)

OpenStudy (dumbsearch2):

See attached problem...

Parth (parthkohli):

Because \(\rm revenue - profit = cost\), ...

OpenStudy (dumbsearch2):

yeeah?

Parth (parthkohli):

You have to calculate \(\rm revenue - profit\) to get the cost.

OpenStudy (dumbsearch2):

like? what? which formula how?

Parth (parthkohli):

``` var x; revenue = 200 - 2x^2; profit = 3x^2 + 245x + 210; cost = revenue - profit; print(cost);

Parth (parthkohli):

I don't know what programming language that is, but do you get what this question means looking at the program?

Parth (parthkohli):

Correction: ``` print(cost,x); ```

OpenStudy (dumbsearch2):

Yeah! :) Wow the way you declare the variables are like JavaScript, and it's printed kinda like PHP lol. :) But it's so funny now that you explain it that way I understand totally! :D

OpenStudy (dumbsearch2):

? cost,x?

OpenStudy (dumbsearch2):

what's the var x?

OpenStudy (dumbsearch2):

undeclared? empty?

Parth (parthkohli):

Okay, I don't wanna confuse you, so take it as `print(cost)`

Parth (parthkohli):

undeclared.

OpenStudy (dumbsearch2):

Hey I get 10 when solving revenue, but I can't seem to be able to solve profit, I run into issues... (you know, this one: 3x^2 + 245x + 210) Thanks soo much for your time! :D

Parth (parthkohli):

In your question, you just assume \(x\) is something and then you find \(200 - 2x^2 - (3x^2 +245x +210)\).

Parth (parthkohli):

Sorry, that is \(200x - 2x^2 -(3x^2 + 245x + 210)\).

OpenStudy (dumbsearch2):

Thank so much for your response. :) I will try to solve it further tomorrow. Thanks. :D

OpenStudy (dumbsearch2):

morning

Parth (parthkohli):

It's pretty easy, actually. :-)

OpenStudy (dumbsearch2):

Thank you, it worked. :)